a swimming pool can be filled by 3 water pipes, A, B and C. Pipe C alone can fill the pool in 8 hours. If pipes B and C are used together, it can be filled in 6 hours. If pipes A and B are used together, it takes 10 hours. How long does it take to fill the pool if all three pipes are used?
Answer by lwsshak3(11628)

a swimming pool can be filled by 3 water pipes, A, B and C. Pipe C alone can fill the pool in 8 hours. If pipes B and C are used together, it can be filled in 6 hours. If pipes A and B are used together, it takes 10 hours. How long does it take to fill the pool if all three pipes are used?
***
1/8=fill rate of C
1/x=fill rate of B
1/y=fill rate of A
1/z=fill rate when all three pipes are used
sum of individual rates=rate working together
...
1/8+1/x=1/6
1/x=1/6-1/8=4/24-3/24=1/24
1/x=1/24(fill rate of B)
..
1/x+1/y=1/10
1/24+1/y=1/10
LCD:240
1/y=1/10-1/24=(24-10)/240=14/240=7/120(fill rate of A)
..
1/8+1/24+7/120=1/z
LCD:120
(15+5+7)/120=27/120=9/40
How long does it take to fill the pool if all three pipes are used? 40/9 hrs